OpenStudy (anonymous):

Simplify each of the following. Write your answer using only positive exponents. Assume that all variables represent nonzero numbers. x^3/x^-7 Please explain how to do this.

OpenStudy (anonymous):

okay do you how to convert a negative exponent to a positive one?

OpenStudy (anonymous):

Do you flip the equation?

OpenStudy (skullpatrol):

$$\frac{x^3}{x^{-7}}$$

OpenStudy (anonymous):

kinda \[x^{-n}=\frac{ 1 }{ x^n }\]

OpenStudy (anonymous):

So then what? Just add the exponents 3 & 7?

OpenStudy (anonymous):

yup you got it.

OpenStudy (anonymous):

Okay thanks!

OpenStudy (anonymous):

also you could have tried subtracting the exponets instead of converting them.\[\frac{ x^n }{ x^k }=x ^{n-k}\]